OpenStudy (astrophysics):

A particle of mass m slides down an inclined plane under the influence of gravity. If the motion is resisted by a force \(f = kmv^2\), show that the time required to move a distance d after starting from rest is \(t = \frac{ \cosh^{-1}(e^{kd}) }{ \sqrt{kg \sin \theta} }\) where \(\theta\) is the anlge of inclination of the plane.

ganeshie8 (ganeshie8):

\[m\ddot{x} = mg\cos\theta - km {\dot{x}}^2\]

OpenStudy (astrophysics):

cos?

ganeshie8 (ganeshie8):

\[m\ddot{x} = mg\sin\theta - km {\dot{x}}^2\]

OpenStudy (astrophysics):

Too easy for you :p

ganeshie8 (ganeshie8):

How are we gonna solve it ?

OpenStudy (astrophysics):

Try rewriting it in a different form

ganeshie8 (ganeshie8):

lumping the constants, i get \[\ddot{x} + k\dot{x}^2 = l\] \(k, l\) are constants

OpenStudy (astrophysics):

Well I sort of did it a messy way I guess, got a separable, if you write it as \[\frac{ 1 }{ k } \frac{ dv }{ \frac{ g }{ k }\sin \theta -v^2 } = dt\]

OpenStudy (astrophysics):

Yeah looking up the integral is probably best I think

ganeshie8 (ganeshie8):

Ahh reduction of order !

OpenStudy (dan815):

since we simply have to show it after writing the DE , showing that equation satisfies the ODE is good enough

OpenStudy (dan815):

so rewrite d=f(t) and show that the ODE is satisfied

ganeshie8 (ganeshie8):

yeahh that will do also solving it isn't that hard... only the algebra is a pain..